Find the number of ways of selecting 9 balls from 6 red balls, 5 white balls and 4 blue balls if each selection consists of 3 balls of each color
1
Expert's answer
2020-03-05T09:41:04-0500

Number of ways of selecting 3 red balls from 6 red balls:


"\frac{6!}{3!(6-3)!}=20"

Number of ways of selecting 3 white balls from 5 white balls:


"\frac{5!}{3!(5-3)!}=10"

Number of ways of selecting 3 blue balls from 4 blue balls:


"\frac{4!}{3!(4-3)!}=4"

So,


"N=(20)(10)(4)=800"